# Hopf Bifurcation of Nonlinear Non-Hermitian Skin Effect

### v1

@@ Acknowledgements.
K.K. thanks Naomichi Hatano and Tomi Ohtsuki for helpful discussion. K.K. thanks Kazuki Yamamoto for letting him know about StreamPlot of Mathematica. K.K. is supported by MEXT KAKENHI Grant-in-Aid for Transformative Research Areas A “Extreme Universe” No. 24H00945. D.N. is supported by JSPS KAKENHI Grant No. 24K22857.
@@
I Existence and uniqueness of limit cycle
@@
$\frac{d^{2}\psi}{dx^{2}}+f\left(\psi\right)\frac{d\psi}{dx}+g\left(\psi\right)=0$ (I.1)
@@
$f\left(\psi\right)\coloneqq 2\left(\varepsilon\psi^{2}-\gamma\right),\quad g\left(\psi\right)\coloneqq 2E\psi.$ (I.2)
@@
[…]
@@
II Singular perturbation theory for weak nonlinearity
II.1 Continuum model
$\tilde{\gamma}\coloneqq\frac{\gamma}{\varepsilon}=\mathcal{O}\left(1\right).$ (II.1)
@@ II.1 Continuum model
$\partial_{x}^{2}\psi+2\varepsilon\left(\psi^{2}-\tilde{\gamma}\right)\partial_{x}\psi+2E\psi=0.$ (II.2)
$\psi=\psi_{0}+\varepsilon\psi_{1}+\mathcal{O}\left(\varepsilon^{2}\right)$ (II.3)
and introduce the two different length scales
$x_{0}\coloneqq x,\quad x_{1}\coloneqq\varepsilon x.$ (II.4)
Given that
$\partial_{x}=\left(\partial_{x}x_{0}\right)\partial_{x_{0}}+\left(\partial_{x}x_{1}\right)\partial_{x_{1}}=\partial_{x_{0}}+\varepsilon\partial_{x_{1}}\eqqcolon\partial_{0}+\varepsilon\partial_{1},$ (II.5)
we have
$\partial_{x}\psi$ $=\left(\partial_{0}+\varepsilon\partial_{1}\right)\left(\psi_{0}+\varepsilon\psi_{1}\right)\simeq\partial_{0}\psi_{0}+\varepsilon\left(\partial_{1}\psi_{0}+\partial_{0}\psi_{1}\right),$ (II.6)
$\partial_{x}^{2}\psi$ $=\left(\partial_{0}^{2}+2\varepsilon\partial_{0}\partial_{1}\right)\left(\psi_{0}+\varepsilon\psi_{1}\right)\simeq\partial_{0}^{2}\psi_{0}+\varepsilon\left(2\partial_{0}\partial_{1}\psi_{0}+\partial_{0}^{2}\psi_{1}\right).$ (II.7)
Since Eq. (II.2) must hold at $\mathcal{O}\left(1\right)$ and $\mathcal{O}\left(\varepsilon\right)$ independently, we obtain
$\partial_{0}^{2}\psi_{0}+2E\psi_{0}=0,$ (II.8)
$\partial_{0}^{2}\psi_{1}+2E\psi_{1}=-2\partial_{0}\partial_{1}\psi_{0}-2\left(\psi_{0}^{2}-\tilde{\gamma}\right)\partial_{0}\psi_{0}.$ (II.9)
The general solution to Eq. (II.8) is given by
$\psi_{0}=r_{0}\sin\theta_{0},\quad\theta_{0}\coloneqq\sqrt{2E}x_{0}+\phi_{0},$ (II.10)
where $r_{0}\geq 0$ and $\phi_{0}\in\left[0,2\pi\right)$ are functions of $x_{1}$ determined by the boundary conditions. Substituting this solution into Eq. (II.9), we obtain
$\partial_{0}^{2}\psi_{1}+2E\psi_{1}=-2\sqrt{2E}\left(\left(\partial_{1}r_{0}\right)\cos\theta_{0}-r_{0}\left(\partial_{1}\phi_{0}\right)\sin\theta_{0}\right)-2\sqrt{2E}r_{0}\left(r^{2}_{0}\sin^{2}\theta_{0}-\tilde{\gamma}\right)\cos\theta_{0}.$ (II.11)
From the formula
$\sin^{2}\theta_{0}\cos\theta_{0}=\frac{1}{4}\left(\cos\theta_{0}-\cos 3\theta_{0}\right),$ (II.12)
the conditions for the elimination of resonance terms yield
$\partial_{1}r_{0}=-r_{0}\left(\frac{r_{0}^{2}}{4}-\tilde{\gamma}\right),$ (II.13)
$r_{0}\left(\partial_{1}\phi_{0}\right)=0.$ (II.14)
From Eq. (II.13), the existence of a limit cycle is dictated by the sign of $\gamma$:
  • For $\gamma<0$, the fixed point $r_{0}=0$ is unique and stable. As a result, we have $r_{0}\to 0$ for $t\to\infty$, implying the absence of limit cycles.
  • For $\gamma>0$, the fixed point $r_{0}=0$ becomes unstable, giving way to the stable fixed point $r_{0}=2\sqrt{\tilde{\gamma}}$. As a result, we have $r_{0}\to 2\sqrt{\tilde{\gamma}}$ for $t\to\infty$, implying the emergence of a limit cycle.
Moreover, Eq. (II.14) indicates that the limit cycle satisfies $\partial_{1}\phi_{0}=0$, further implying that $\phi_{0}$ remains a constant with respect to $x_{1}$. Therefore, the limit cycle for $\gamma>0$, corresponding to a delocalized eigenstate in real space, is described by
$x\simeq 2\sqrt{\frac{\gamma}{\varepsilon}}\sin\left(kx+\mathcal{O}\left(\varepsilon^{2}\right)\right),\quad k\coloneqq\sqrt{2E}$ (II.15)
for $x\to\infty$. Equation (II.9) further reduces to
$\partial_{0}^{2}\psi_{1}+2E\psi_{1}=\frac{\sqrt{2E}}{2}r_{0}^{3}\cos 3\theta_{0}=4\tilde{\gamma}^{3/2}\sqrt{2E}\cos 3\theta_{0},$ (II.16)
which admits the solution
$\psi_{1}=r_{1}\sin\theta_{1}-\frac{\tilde{\gamma}^{3/2}}{\sqrt{8E}}\cos 3\theta_{0},\quad\theta_{1}\coloneqq\sqrt{2E}x_{0}+\phi_{1}$ (II.17)
with $r_{1}\geq 0$ and $\phi_{1}\in\left[0,2\pi\right)$ dependent on $x_{1}$.
$\star$
To capture the $\mathcal{O}\left(\varepsilon^{2}\right)$ correction of the energy dispersion, it is more effective to employ the Poincaré-Lindstedt method. We reformulate Eq. (II.2) and obtain
$k^{2}\partial_{z}^{2}\psi+2k\varepsilon\left(\psi^{2}-\tilde{\gamma}\right)\partial_{z}\psi+2E\psi=0,\quad z\coloneqq kx.$ (II.18)
@@
$\psi=\psi_{0}+\varepsilon\psi_{1}+\varepsilon^{2}\psi_{2}+\mathcal{O}\left(\varepsilon^{3}\right),\quad k=\sqrt{2E}+\varepsilon k_{1}+\varepsilon^{2}k_{2}+\mathcal{O}\left(\varepsilon^{3}\right),$ (II.19)
@@
$\partial_{z}^{2}\psi_{0}+\psi_{0}=0,$ (II.20)
$\partial_{z}^{2}\psi_{1}+\psi_{1}=-\sqrt{\frac{2}{E}}\left(k_{1}\left(\partial_{z}^{2}\psi_{0}\right)+\left(\psi_{0}^{2}-\tilde{\gamma}\right)\partial_{z}\psi_{0}\right).$ (II.21)
The solution to Eq. (II.20) with the initial condition $\psi_{0}\left(z=0\right)=0$ is given by
$\psi_{0}=r\sin z\quad\left(r\in\mathbb{R}\right).$ (II.22)
Substituting this into Eq. (II.21), we further have
$\partial_{z}^{2}\psi_{1}+\psi_{1}=-\sqrt{\frac{2}{E}}\left(-k_{1}r\sin z+r\left(r^{2}\sin^{2}z-\tilde{\gamma}\right)\cos z\right).$ (II.23)
@@
$k_{1}r=0,\quad r\left(\frac{r^{2}}{4}-\tilde{\gamma}\right)=0,$ (II.24)
@@
$k_{1}=0,\quad r=2\sqrt{\tilde{\gamma}}.$ (II.25)
Then, Eq. (II.23) is simplified to
$\partial_{z}^{2}\psi_{1}+\psi_{1}=\frac{r^{3}}{2\sqrt{2E}}\cos 3z=\frac{4\tilde{\gamma}^{3/2}}{\sqrt{2E}}\cos 3z,$ (II.26)
@@
$\psi_{1}=c\sin z+\frac{\tilde{\gamma}^{3/2}}{2\sqrt{2E}}\left(\cos z-\cos 3z\right)\quad\left(c\in\mathbb{R}\right).$ (II.27)
Now, the nonlinear Schrödinger equation (II.18) for $\mathcal{O}\left(\varepsilon^{2}\right)$ reads
$\partial_{z}^{2}\psi_{2}+\psi_{2}=-\sqrt{\frac{2}{E}}\left(k_{2}\partial_{z}^{2}\psi_{0}+2\psi_{0}\psi_{1}\partial_{z}\psi_{0}+\left(\psi_{0}^{2}-\tilde{\gamma}\right)\partial_{z}\psi_{1}\right).$ (II.28)
@@
$c=0,\quad k_{2}=-\frac{\tilde{\gamma}^{2}}{4\sqrt{2E}}.$ (II.29)
@@
$\simeq 2\sqrt{\tilde{\gamma}}\sin\left(\left(\sqrt{2E}-\frac{\tilde{\gamma}^{2}}{4\sqrt{2E}}\varepsilon^{2}\right)x\right)+\frac{\varepsilon\tilde{\gamma}^{3/2}}{2\sqrt{2E}}\left(\cos\sqrt{2E}x-\cos 3\sqrt{2E}x\right).$ (II.30)
@@
$k=\sqrt{2E}-\frac{\tilde{\gamma}^{2}}{4\sqrt{2E}}\varepsilon^{2}+\mathcal{O}\left(\varepsilon^{3}\right),\quad\mathrm{i.e.},\quad E=\frac{k^{2}}{2}+\frac{\tilde{\gamma}^{2}}{4}\varepsilon^{2}+\mathcal{O}\left(\varepsilon^{3}\right),$ (II.31)

### v2

@@ Acknowledgements.
We thank Zhao-Fan Cai, Naomichi Hatano, Tomi Ohtsuki, and Zhenyu Xiao for helpful discussion. K.K. thanks Kazuki Yamamoto for letting him know about StreamPlot of Mathematica. K.K. is supported by MEXT KAKENHI Grant-in-Aid for Transformative Research Areas A “Extreme Universe” No. JP24H00945. D.N. is supported by JSPS KAKENHI Grant No. JP24K22857.
Appendix A Singular perturbation theory for weak nonlinearity
$\tilde{\gamma}\coloneqq\frac{\gamma}{\varepsilon}=\mathcal{O}\left(1\right).$ (16)
Under this assumption, the nonlinear Schrödinger equation reads
$\partial_{x}^{2}\psi+2\varepsilon\left(\psi^{2}-\tilde{\gamma}\right)\partial_{x}\psi+2E\psi=0.$ (17)
$\psi=\psi_{0}+\varepsilon\psi_{1}+\mathcal{O}\left(\varepsilon^{2}\right)$ (18)
and introduce the two different length scales
$x_{0}\coloneqq x,\quad x_{1}\coloneqq\varepsilon x.$ (19)
Given that
$\partial_{x}=\partial_{x_{0}}+\varepsilon\partial_{x_{1}}\eqqcolon\partial_{0}+\varepsilon\partial_{1},$ (20)
we have
$\partial_{x}\psi$ $\simeq\partial_{0}\psi_{0}+\varepsilon\left(\partial_{1}\psi_{0}+\partial_{0}\psi_{1}\right),$ (21)
$\partial_{x}^{2}\psi$ $\simeq\partial_{0}^{2}\psi_{0}+\varepsilon\left(2\partial_{0}\partial_{1}\psi_{0}+\partial_{0}^{2}\psi_{1}\right).$ (22)
Since Eq. (17) must hold at $\mathcal{O}\left(1\right)$ and $\mathcal{O}\left(\varepsilon\right)$ independently, we obtain
$\partial_{0}^{2}\psi_{0}+2E\psi_{0}=0,$ (23)
$\partial_{0}^{2}\psi_{1}+2E\psi_{1}=-2\partial_{0}\partial_{1}\psi_{0}-2\left(\psi_{0}^{2}-\tilde{\gamma}\right)\partial_{0}\psi_{0}.$ (24)
The general solution to Eq. (23) is given by
$\psi_{0}=r_{0}\sin\theta_{0},\quad\theta_{0}\coloneqq\sqrt{2E}x_{0}+\phi_{0},$ (25)
where $r_{0}\geq 0$ and $\phi_{0}\in\left[0,2\pi\right)$ are functions of $x_{1}$ determined by the boundary conditions. Substituting this solution into Eq. (24), we obtain
$\partial_{0}^{2}\psi_{1}+2E\psi_{1}=-2\sqrt{2E}\left(\left(\partial_{1}r_{0}\right)\cos\theta_{0}-r_{0}\left(\partial_{1}\phi_{0}\right)\sin\theta_{0}\right)$
$\qquad\qquad\qquad-2\sqrt{2E}r_{0}\left(r^{2}_{0}\sin^{2}\theta_{0}-\tilde{\gamma}\right)\cos\theta_{0}.$ (26)
From the formula
$\sin^{2}\theta_{0}\cos\theta_{0}=\frac{1}{4}\left(\cos\theta_{0}-\cos 3\theta_{0}\right),$ (27)
the conditions for the elimination of resonance terms yield
[…]
$r_{0}\left(\partial_{1}\phi_{0}\right)=0.$ (29)
From Eq. (28), the existence of a limit cycle is dictated by the sign of $\gamma$:
  • For $\gamma<0$, the fixed point $r_{0}=0$ is unique and stable. As a result, we have $r_{0}\to 0$ for $x\to\infty$, implying the absence of limit cycles.
  • For $\gamma>0$, the fixed point $r_{0}=0$ becomes unstable, giving way to the stable fixed point $r_{0}=2\sqrt{\tilde{\gamma}}$. As a result, we have $r_{0}\to 2\sqrt{\tilde{\gamma}}$ for $x\to\infty$, implying the emergence of a limit cycle.
Moreover, Eq. (29) indicates that the limit cycle satisfies $\partial_{1}\phi_{0}=0$, further implying that $\phi_{0}$ remains a constant with respect to $x_{1}$. Therefore, the limit cycle for $\gamma>0$, corresponding to a delocalized eigenstate in real space, is described by
$\psi_{0}\simeq 2\sqrt{\frac{\gamma}{\varepsilon}}\sin\left(kx+\mathcal{O}\left(\varepsilon^{2}\right)\right),\quad k\coloneqq\sqrt{2E}$ (30)
for $x\to\infty$. Equation (24) further reduces to
$\partial_{0}^{2}\psi_{1}+2E\psi_{1}=4\tilde{\gamma}^{3/2}\sqrt{2E}\cos 3\theta_{0},$ (31)
which admits the solution
$\psi_{1}=r_{1}\sin\theta_{1}-\frac{\tilde{\gamma}^{3/2}}{\sqrt{8E}}\cos 3\theta_{0},\quad\theta_{1}\coloneqq\sqrt{2E}x_{0}+\phi_{1}$ (32)
with $r_{1}\geq 0$ and $\phi_{1}\in\left[0,2\pi\right)$ dependent on $x_{1}$. In the Supplemental Material [123], we further derive the $\mathcal{O}\left(\varepsilon^{2}\right)$ correction to the energy dispersion using the Poincaré-Lindstedt method.
In passing, the leading-order wave function for $\gamma<0$ is obtained from Eq. (28) as
$\psi_{0}\simeq 2\sqrt{-\frac{\gamma}{\varepsilon}}\frac{e^{\gamma\left(x-4c\right)}}{\sqrt{1-e^{2\gamma\left(x-4c\right)}}}\sin kx,\quad k\coloneqq\sqrt{2E},$ (33)
where the constant $c$ is determined by the boundary conditions. This shows the nontrivial correction of skin states induced by nonlinearity. Specifically, $\psi_{0}$ contains not only original skin states $e^{\gamma x}$ but also additional decaying contributions $e^{\left(2n+1\right)\gamma x}$ ($n\in\mathbb{N}_{>0}$), which constitutes a signature of the enhanced localization due to nonlinearity.
Appendix B Limit cycle for strong nonlinearity
We analyze a delocalized eigenstate associated with a limit cycle in phase space under the regime of strong nonlinearity $\varepsilon\gg 1$. The nonlinear Schrödinger equation for the nonlinear Hatano-Nelson model reads
$\partial_{x}\psi$ $=\varepsilon\left(\chi-2F\left(\psi\right)\right),$ (34)
$\partial_{x}\chi$ $=-\frac{2E}{\varepsilon}\psi,$ (35)
with
$F\left(\psi\right)\coloneqq\frac{1}{3}\psi^{3}-\frac{\gamma}{\varepsilon}\psi,\quad\chi\coloneqq\frac{1}{\varepsilon}\partial_{x}\psi+2F\left(\psi\right).$ (36)
Owing to $\varepsilon\gg 1$, Eq. (34) indicates that an arbitrary initial point in phase space $\left(\psi,\chi\right)$ rapidly converges to the nullcline $\chi=2F\left(\psi\right)$. Subsequently, it slowly drifts along this nullcline according to Eq. (35), ultimately leading to the convergence at the fixed point $\left(0,0\right)$ for $\gamma\leq 0$ and the emergence of a limit cycle for $\gamma>0$. Since the cubic function $\chi=2F\left(\psi\right)$ attains extrema $\pm\chi_{0}\coloneqq\pm\left(4/3\right)\left(\gamma/\varepsilon\right)^{3/2}$ at $\psi=\pm\psi_{0}=\pm\sqrt{\gamma/\varepsilon}$, the period of the limit cycle, corresponding to the wavelength of the delocalized eigenstate, is evaluated as
$\lambda\simeq 2\int_{2\psi_{0}}^{\psi_{0}}\left(-\frac{\varepsilon}{2E\psi}\right)\left(2F^{\prime}\left(\psi\right)\right)d\psi=\left(3-2\log 2\right)\frac{\gamma}{E}.$ (37)
Defining a wave number as $k\coloneqq 2\pi/\lambda$, we have the energy dispersion relation in Eq. (8).
Figure 4: Dynamics of the nonlinear Hatano-Nelson model on the lattice under the open boundary conditions ($L=49$). Plots of $\left|\psi_{n}\left(t\right)\right|$ as functions of site $n$ and time $t$ for (a) $\gamma=0$, $\varepsilon=0$, (b) $\gamma=0.1$, $\varepsilon=0$, (c) $\gamma=-0.1$, $\varepsilon=0$, and (d) $\gamma=0.1$, $\varepsilon=0.5$. The initial conditions are set as $\psi_{n}\left(t=0\right)=\delta_{n,(L+1)/2}$.
Appendix C Nonlinear dynamics
[…]
@@ Appendix C Nonlinear dynamics
Appendix A Existence and uniqueness of limit cycle
@@
$\frac{d^{2}\psi}{dx^{2}}+f\left(\psi\right)\frac{d\psi}{dx}+g\left(\psi\right)=0$ (A.1)
@@
$f\left(\psi\right)\coloneqq 2\left(\varepsilon\psi^{2}-\gamma\right),\quad g\left(\psi\right)\coloneqq 2E\psi.$ (A.2)
@@
    $F\left(\psi\right)\coloneqq\int_{0}^{\psi}f\left(\chi\right)d\chi=2\left(\frac{\varepsilon}{3}\psi^{3}-\gamma\psi\right).$ (A.3)
@@
Appendix B Singular perturbation theory for weak nonlinearity
B.1 Continuum model
We perturbatively derive a delocalized eigenstate associated with a limit cycle in phase space for weak nonlinearity $0\leq\varepsilon\ll 1$. We focus on the vicinity of the Hopf bifurcation and assume that the linear non-Hermitian parameter $\gamma$ is of the same order as the nonlinear non-Hermitian parameter $\varepsilon$, i.e.,
$\tilde{\gamma}\coloneqq\frac{\gamma}{\varepsilon}=\mathcal{O}\left(1\right).$ (B.1)
@@ B.1 Continuum model
$\partial_{x}^{2}\psi+2\varepsilon\left(\psi^{2}-\tilde{\gamma}\right)\partial_{x}\psi+2E\psi=0.$ (B.2)
To capture the $\mathcal{O}\left(\varepsilon^{2}\right)$ correction of the energy dispersion, we here employ the Poincaré-Lindstedt method. We reformulate Eq. (B.2) and obtain
$k^{2}\partial_{z}^{2}\psi+2k\varepsilon\left(\psi^{2}-\tilde{\gamma}\right)\partial_{z}\psi+2E\psi=0,\quad z\coloneqq kx.$ (B.3)
@@
$\psi=\psi_{0}+\varepsilon\psi_{1}+\varepsilon^{2}\psi_{2}+\mathcal{O}\left(\varepsilon^{3}\right),\quad k=\sqrt{2E}+\varepsilon k_{1}+\varepsilon^{2}k_{2}+\mathcal{O}\left(\varepsilon^{3}\right),$ (B.4)
@@
$\partial_{z}^{2}\psi_{0}+\psi_{0}=0,$ (B.5)
$\partial_{z}^{2}\psi_{1}+\psi_{1}=-\sqrt{\frac{2}{E}}\left(k_{1}\left(\partial_{z}^{2}\psi_{0}\right)+\left(\psi_{0}^{2}-\tilde{\gamma}\right)\partial_{z}\psi_{0}\right).$ (B.6)
The solution to Eq. (B.5) with the initial condition $\psi_{0}\left(z=0\right)=0$ is given by
$\psi_{0}=r\sin z\quad\left(r\in\mathbb{R}\right).$ (B.7)
Substituting this into Eq. (B.6), we further have
$\partial_{z}^{2}\psi_{1}+\psi_{1}=-\sqrt{\frac{2}{E}}\left(-k_{1}r\sin z+r\left(r^{2}\sin^{2}z-\tilde{\gamma}\right)\cos z\right).$ (B.8)
@@
$k_{1}r=0,\quad r\left(\frac{r^{2}}{4}-\tilde{\gamma}\right)=0,$ (B.9)
@@
$k_{1}=0,\quad r=2\sqrt{\tilde{\gamma}}$. (B.10)
Then, Eq. (B.8) is simplified to
$\partial_{z}^{2}\psi_{1}+\psi_{1}=\frac{r^{3}}{2\sqrt{2E}}\cos 3z=\frac{4\tilde{\gamma}^{3/2}}{\sqrt{2E}}\cos 3z,$ (B.11)
@@
$\psi_{1}=c\sin z+\frac{\tilde{\gamma}^{3/2}}{2\sqrt{2E}}\left(\cos z-\cos 3z\right)\quad\left(c\in\mathbb{R}\right).$ (B.12)
Now, the nonlinear Schrödinger equation (B.3) for $\mathcal{O}\left(\varepsilon^{2}\right)$ reads
$\partial_{z}^{2}\psi_{2}+\psi_{2}=-\sqrt{\frac{2}{E}}\left(k_{2}\partial_{z}^{2}\psi_{0}+2\psi_{0}\psi_{1}\partial_{z}\psi_{0}+\left(\psi_{0}^{2}-\tilde{\gamma}\right)\partial_{z}\psi_{1}\right).$ (B.13)
@@
$c=0,\quad k_{2}=-\frac{\tilde{\gamma}^{2}}{4\sqrt{2E}}.$ (B.14)
@@
$\simeq 2\sqrt{\tilde{\gamma}}\sin\left(\left(\sqrt{2E}-\frac{\tilde{\gamma}^{2}}{4\sqrt{2E}}\varepsilon^{2}\right)x\right)+\frac{\varepsilon\tilde{\gamma}^{3/2}}{2\sqrt{2E}}\left(\cos\sqrt{2E}x-\cos 3\sqrt{2E}x\right).$ (B.15)
@@
$k=\sqrt{2E}-\frac{\tilde{\gamma}^{2}}{4\sqrt{2E}}\varepsilon^{2}+\mathcal{O}\left(\varepsilon^{3}\right),\quad\mathrm{i.e.},\quad E=\frac{k^{2}}{2}+\frac{\tilde{\gamma}^{2}}{4}\varepsilon^{2}+\mathcal{O}\left(\varepsilon^{3}\right),$ (B.16)